OpenStudy (anonymous):

Solve the logarithmic equation below:

OpenStudy (anonymous):

\[7-\log_{3}8x=2 \]

OpenStudy (anonymous):

7 minus log base of 3 multiplied by 8x = 2

OpenStudy (anonymous):

\[7 - \log _{2}8x = 2\rightarrow -\log _{2}8x = -5\rightarrow \frac{ 1 }{ 2^{5} } = 8x \rightarrow x = \frac{ 1 }{ 256 }\]

OpenStudy (anonymous):

That's wrong but thanks. I already found out the answer.

OpenStudy (anonymous):

ok, could you tell me which part was incorrect?

OpenStudy (anonymous):

You used log base of 2 instead of 3

OpenStudy (anonymous):

ah, ok three and 2 look to similar in this font.

OpenStudy (anonymous):

Yea I know I don't blame you, thanks again though.

OpenStudy (anonymous):

no problem.